OpenStudy (anonymous):

[10.01] Which factor can you cancel from the numerator and denominator to simplify 12 minus 6 times x all over 5 times x minus 10.? [Hint: (a−x) = − (x − a)] 6 over 5. 5 over 6. x + 2 x − 2

OpenStudy (anonymous):

\[\frac{ 12 - 6x }{ 5x- 10 } = \frac{ 6(2 - x) }{ 5(x-2) } = \frac{ -6(x-2) }{ 5(x-2) }\] i used the hint for the last step. answer is D

OpenStudy (anonymous):

can you help me on this one too [10.01] Simplify x squared minus 36 all over x squared minus 12 times x plus 36.. x minus 6 all over x minus 3. x plus 6 all over x minus 6. x minus 3 all over x plus 3. x plus 3 all over x plus 6.